# Spatial-Aware Decision-Making with Ring Attractors in Reinforcement Learning Systems

### v2

@@ A.4.1 Neutral Inhibitory Action Implementation
For example, in games like Seaquest or Asterix, this central positioning means the ’no-op’ action has consistent, strong connections to all directional actions. This arrangement preserves the spatial relationships between directional actions while ensuring the neutral action remains equally accessible from any game state. The constant distance of 1 to all other actions makes transitioning to or from the neutral action as natural as moving between adjacent directional actions in the ring.
A.4.2 Deep Learning Double Ring Attractor Equations
For a double ring configuration in our DL implementation as presented in thee experiments, Section 4.2, the weighted connections are defined as follows:
@@ A.4.2 Deep Learning Double Ring Attractor Equations
$[\mathbf{V}_{ii}]_{m,n}=\frac{1}{\tau}\Phi_{\theta}(s)^{T}e^{d(m,n)/\lambda},$ (15)
@@
$Q(s,a)=\beta\tanh\left(\begin{bmatrix}\frac{1}{\tau}\Phi_{\theta}(s_{t})^{T}\mathbf{V}_{11}+h_{t-1}(v)^{T}\mathbf{U}_{11}&\frac{\kappa}{\tau}\Phi_{\theta}(s_{t})^{T}\mathbf{V}_{12}+\kappa h_{t-1}(v)^{T}\mathbf{U}_{12}\\ \frac{\kappa}{\tau}\Phi_{\theta}(s_{t})^{T}\mathbf{V}_{21}+\kappa h_{t-1}(v)^{T}\mathbf{U}_{21}&\frac{1}{\tau}\Phi_{\theta}(s_{t})^{T}\mathbf{V}_{22}+h_{t-1}(v)^{T}\mathbf{U}_{22}\end{bmatrix}\right),$ (18)
@@
The cross-coupling matrices ($\kappa\mathbf{V}{12}$, $\kappa\mathbf{V}{21}$, $\kappa\mathbf{U}{12}$, and $\kappa\mathbf{U}{21}$) maintain a circular topology similar to the individual rings. A neuron at a particular position in the first ring connects most strongly to the neuron at the corresponding position in the second ring, with connection strength decreasing based on circular distance. This structured cross-coupling preserves spatial alignment between the two action dimensions while allowing semi-independent operation through the learnable coupling factor $\kappa$.
The final output provides action-values for both action dimensions simultaneously, preserving the spatial relationships within each ring while allowing for weak coupling between the rings.
A.4.3 Extension to N Ring Configurations
The double ring implementation extends to $R$ rings through a block matrix structure, where each ring encodes a distinct action dimension. For $R$ rings, the architecture uses block matrices $V_{\text{multi}}\in\mathbb{R}^{RN\times RM}$ and $U_{\text{multi}}\in\mathbb{R}^{RN\times RN}$, with diagonal blocks preserving individual ring dynamics and off-diagonal blocks handling cross-ring interactions via coupling parameter $\kappa$, as seen in Section A.4.2. While computational complexity scales as $O(R^{2})$, selective coupling between only related dimensions creates a sparse structure with effective $O(R)$ complexity. This makes the approach viable for complex action spaces where actions decompose into multiple semi-independent planes, such as games combining movement, combat, and resource management dimensions.
A.5 Models and Environments Implementation
We provide implementation details for both our models and the tested environments. For model implementations, EffZeroRA was applied across the Atari benchmark suite, while BDQNRA-UA and DDQNRA were specifically implemented for Highway and Mario Bros. Table 2 details the configuration of action spaces and ring architectures for each environment. The environments required different ring configurations based on their control schemes, ranging from single-ring implementations for basic movement to double-ring setups for more complex action spaces that combine movement and specialised actions. Each ring’s topology was designed to preserve the natural relationships between actions, with central inhibitory actions included where appropriate as “no action”.

### v3

@@ A.7.1 Neutral Inhibitory Action Implementation
For example, in games like Seaquest or Asterix, this central positioning means the “no-op” action has consistent, strong connections to all directional actions. This arrangement preserves the spatial relationships between directional actions while ensuring the neutral action remains equally accessible from any game state. The constant distance of 1 to all other actions makes transitioning to or from the neutral action as natural as moving between adjacent directional actions in the ring.
A.7.2 Deep Learning Double Ring Attractor Equations
For a double ring configuration in our DL implementation as presented in the experiments, Section 4.2, the weighted connections are defined as follows:
@@ A.7.2 Deep Learning Double Ring Attractor Equations
$\mathbf{V}_{ii}]_{m,n}=\frac{1}{\tau}\Phi_{\theta}(s)^{T}e^{-d(m,n)/\lambda},$ (27)
@@
[…]
@@
The cross-coupling matrices ($\kappa\mathbf{V}{12}$, $\kappa\mathbf{V}{21}$, $\kappa\mathbf{U}{12}$, and $\kappa\mathbf{U}{21}$) maintain a circular topology similar to the individual rings. A neuron at a particular position in the first ring connects most strongly to the neuron at the corresponding position in the second ring, with connection strength decreasing based on circular distance. This structured cross-coupling preserves spatial alignment between the two action dimensions while allowing semi-independent operation through the learnable coupling factor $\kappa$. The final output provides action-values for both action dimensions simultaneously, preserving spatial relationships within each ring while allowing weak coupling between them.
A.7.3 Extension to N Ring Configurations
The double ring implementation extends to $R$ rings through a block matrix structure, where each ring encodes a distinct action dimension. For $R$ rings, the architecture uses block matrices $V_{\text{multi}}\in\mathbb{R}^{RN\times RM}$ and $U_{\text{multi}}\in\mathbb{R}^{RN\times RN}$, with diagonal blocks preserving individual ring dynamics and off-diagonal blocks handling cross-ring interactions via coupling parameter $\kappa$, as seen in Section A.7.2. While computational complexity scales as $O(R^{2})$, selective coupling between only related dimensions creates a sparse structure with effective $O(R)$ complexity. This makes the approach viable for complex action spaces where actions decompose into multiple semi-independent planes, for example, in games that combine movement, combat, and resource-management dimensions.
A.8 Models and Environments Implementation
@@
We provide implementation details for both our models and the tested environments. For model implementations, EffZeroRA was applied across the Atari benchmark suite, while BDQNRA-UA and DDQNRA were specifically implemented for Highway and Mario Bros. Table 6 details the configuration of action spaces and ring architectures for each environment. The environments required different ring configurations based on their control schemes, ranging from single-ring implementations for basic movement to double-ring setups for more complex action spaces that combine movement and specialized actions. Each ring topology was designed to preserve the natural relationships between actions, with central inhibitory actions included where appropriate as “no action”.
A.8.1 Computational Resources
For the Atari 100K benchmark [5] experiments, we utilized a high-performance computing cluster equipped with 10 NVIDIA A100 GPUs (each with 80 GB memory), 512 GB of RAM, and 128 Intel Xeon CPU cores running at 2.4 GHz. The cluster environment enabled parallel execution of experiments across multiple seeds and games simultaneously, with training times ranging from 8 to 10 hours per game, depending on complexity.
For the other environments, Highway [21] and Super Mario Bros [15], we employed a local workstation with an Intel Xeon processor (28 cores, 2.1 GHz), 125 GB RAM, and dual NVIDIA RTX A5000 GPUs (total 48 GB combined memory). Training on the local machine typically required 6-12 hours per environment, with the CTRNN implementation incurring additional computational overhead as noted in Section 4.1. All experiments were conducted using PyTorch 1.12 with CUDA 11.6. The total computational cost is estimated at approximately 8,000 GPU-hours across all experiments and ablation studies.